OpenStudy (anonymous):

Write the equation in SLOPE-INTERCEPT FORM for a line that contains (-6,1) (2,-3)  

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

take the slope of (-6,1) and (2,-3). I got -1/2 Then write y=-1/2x+b plug in a point 1=-1/2(-6)+b b=-2 so your equation is y=-1/2x-2

OpenStudy (anonymous):

thank you